Three men, four women and six children can complete a work in seven days. A women does double the work a man does and a child does half the work a man does. How many women alone can complete this work in 7 days.
Answer by lwsshak3(6461)

Three men, four women and six children can complete a work in seven days. A women does double the work a man does and a child does half the work a man does. How many women alone can complete this work in 7 days.
...
let x=days 4 women take to complete the job by themselves
1/x=4 women's work rate
1/2x=3 men's work rate
1/4x=6 children's work rate
work rate when working together=1/7
1/x+1/2x+1/4x=1/7
LCD: 28x
28+14+7=4x
4x=49
x=49/4
4 women working together by themselves can complete the job in 49/4 days
1/u=each women's work rate
4/u=4/49
u=49 days each women would take to complete the job by herself
so 7 women could complete the job in 7 days working together
...